OpenStudy (anonymous):

verify: tanxcos(2x)=sin(2x)-tanx

OpenStudy (abb0t):

HINT: \(\tan(x) = \frac{ \sin(x) }{ \cos(x) }\) and \(\cos(2x)= 2\cos(x)-1\) or \(\cos(2x)= \cos^2(x)-\sin^2(x)\)

OpenStudy (abb0t):

also: \(\cos(2x)=1-2\sin^2(x)\)

OpenStudy (abb0t):

error on the first part, i meant: \(\cos(2x) = 2\cos^2(x)-1\)

OpenStudy (anonymous):

\[\tan x \times \cos 2x = \left( \sin x / \cos x \right) \times (2(\cos x)^{2} -1)\] \[= 2*Cos x *\sin x - \tan x = \sin (2x) - tanx\]

OpenStudy (anonymous):

thanks guys ive literaly been working on that problem for an hour.